the length of a rectangle mailing is four centimeters less than twice the width. the perimeter is 40 centimeters. Find the dimensions of the label?
Answer by mananth(16949)

the length of a rectangle mailing is four centimeters less than twice the width. the perimeter is 40 centimeters. Find the dimensions of the label?
Let the width be x
twice the width = 2x
length is 4cms less than twice the width
so length = 2x-4
Perimeter of rectangle -= 2*( length + width)
= 2 * ( x+ 2x -4)
=2*(3x-4)
= 6x-8
6x-8 = 40
6x= 40+8=48
x=8
2x-4 = width
2*8-4=12
width 8, length 12